# Whisper-Flamingo: Integrating Visual Features into Whisper for Audio-Visual Speech Recognition and Translation

### v2

@@ Abstract
Audio-Visual Speech Recognition (AVSR) uses lip-based video to improve performance in noise. Since videos are harder to obtain than audio, the video training data of AVSR models is usually limited to a few thousand hours. In contrast, speech models such as Whisper are trained with hundreds of thousands of hours of data, and thus learn a better speech-to-text decoder. The huge training data difference motivates us to adapt Whisper to handle video inputs. Inspired by Flamingo which injects visual features into language models, we propose Whisper-Flamingo which integrates visual features into the Whisper speech recognition and translation model with gated cross attention. Our models achieve state-of-the-art ASR WER (0.68%) and AVSR WER (0.76%) on LRS3. Audio-visual Whisper-Flamingo outperforms audio-only Whisper on English speech recognition and En-X translation for 6 languages in noisy conditions. Moreover, Whisper-Flamingo is versatile and conducts all of these tasks using one set of parameters, while prior methods are trained separately on each language.
@@ 1 Introduction
On the English (En) LRS3 video dataset [10], our models achieve State-of-the-Art (SOTA) ASR WER (0.68%) and AVSR WER (0.76%). Our novel audio-visual Whisper-Flamingo significantly outperforms the audio-only Whisper baseline in noise. Moreover, Whisper-Flamingo achieves competitive noise-robust results compared to prior audio-visual models. Next, we demonstrate Whisper’s multilingual capabilities by extending Whisper-Flamingo for En-X translation on the MuAViC dataset [11]. Our model performs En transcription and En-X translation into 6 other languages, while the previous audio-visual SOTA requires fine-tuning on each language separately. Once again, Whisper-Flamingo significantly outperforms audio-only Whisper in noise for both En transcription and En-X translation. Code and models at https://github.com/roudimit/whisper-flamingo

### v3

@@ Abstract
Audio-Visual Speech Recognition (AVSR) uses lip-based video to improve performance in noise. Since videos are harder to obtain than audio, the video training data of AVSR models is usually limited to a few thousand hours. In contrast, speech models such as Whisper are trained with hundreds of thousands of hours of data, and thus learn a better speech-to-text decoder. The huge training data difference motivates us to adapt Whisper to handle video inputs. Inspired by Flamingo which injects visual features into language models, we propose Whisper-Flamingo which integrates visual features into the Whisper speech recognition and translation model with gated cross attention. Our models achieve state-of-the-art ASR WER (0.68%) and AVSR WER (0.76%) on LRS3, and state-of-the-art ASR WER (1.3%) and AVSR WER (1.4%) on LRS2. Audio-visual Whisper-Flamingo outperforms audio-only Whisper on English speech recognition and En-X translation for 6 languages in noisy conditions. Moreover, Whisper-Flamingo is versatile and conducts all of these tasks using one set of parameters, while prior methods are trained separately on each language.
@@ 1 Introduction
On the English (En) LRS3 video dataset [10], our models achieve State-of-the-Art (SOTA) ASR WER (0.68%) and AVSR WER (0.76%). On LRS2 [11], our models achieve SOTA ASR WER (1.3%) and AVSR WER (1.4%). Our novel audio-visual Whisper-Flamingo significantly outperforms the audio-only Whisper baseline in noise. Moreover, Whisper-Flamingo achieves competitive noise-robust results compared to prior audio-visual models. Next, we demonstrate Whisper’s multilingual capabilities by extending Whisper-Flamingo for En-X translation on the MuAViC dataset [12]. Our model performs En transcription and En-X translation into 6 other languages, while the previous audio-visual SOTA requires fine-tuning on each language separately. Once again, Whisper-Flamingo significantly outperforms audio-only Whisper in noise for both En transcription and En-X translation. Code and models at https://github.com/roudimit/whisper-flamingo
@@ 3.4 Whisper-Flamingo En-X Speech Translation
Table 5: Results for English transcription on LRS2. We report results on the original test set (Clean). Hours of unlabeled & labeled audio-visual data used to train each model are shown. Note that we used AV-HuBERT fine-tuned on LRS3 433h.
[…]
Model Unlabeled Labeled ASR AVSR
Audio-Visual methods trained only on transcribed videos
TM-seq2seq [11] - 223 9.7 8.5
CTC/Attention [53] - 380 8.3 7.0
TDNN [54] - 223 6.7 5.9
CM-seq2seq [20] - 380 3.9 3.7
Efficient Conformer [55] - 818 2.4 2.3
AutoAVSR [21] - 3,448 1.5 1.5
Audio-Visual SSL methods
UniAVSR [5] - 380 2.7 2.6
RAVEn [56] 1,759 223 2.5 -
USR [57] 1,759 223 1.9 1.9
Our audio-only Whisper Zero-shot baselines
Whisper-Medium, Beam 1 - - 5.2 -
Whisper-Large, Beam 1 - - 5.5 -
Our models trained with noise
Whisper-Medium FT, Beam 1 - 223 1.3 -
Whisper-Flamingo, Beam 1 1,759 223† - 1.4
4 Experiments on LRS2
We use our framework to conduct experiments on the LRS2 dataset [11]. We followed AutoAVSR [21] to create a 223h training set, 0.6h validation set, and 0.5h test set. We add noise to the audio while training our models. We report the ASR and AVSR WER on the LRS2 test set (clean) in Table 5. Comparing zero-shot Whisper Medium and Whisper Large V2, the medium model performs slightly better (5.2% vs 5.5%), therefore we use Whisper medium for our experiments. Our fine-tuned Whisper and Whisper-Flamingo models have 769M and 1.39B parameters respectively. Compared to zero-shot Whisper, fine-tuned Whisper improves the ASR WER from 5.2% to 1.3%, achieving a new SOTA ASR WER on LRS2. Finally, Whisper-Flamingo achieves SOTA AVSR WER of 1.4%. Whisper-Flamingo outperforms recent SSL method USR [57] (1.9%) trained with the same hours of video and outperforms AutoAVSR [21] (1.5%) trained with more hours of video. Finally, Whisper-Flamingo outperforms Whisper in noisy conditions (shown in the Appendix, Section 7.3).